# Gregorio Curello Mannheim Ludvig Sinander

### v1

@@ 6 Dominance by objects (rather than mixtures)
U(row,column) ij
Note that $y\in\text{NWD}(X,\mathcal{S},U)$, and the mixture $p\in\Delta(X)$ given by $p(x)\coloneqq\alpha$, $p(y)\coloneqq 0$ and $p(z)\coloneqq 1-\alpha$ satisfies $\sum_{w\in X}\Phi(U(w,\cdot))p(w)>\Phi(U(y,\cdot))$, so that $y\notin\text{NWD}(X,\mathcal{S},\Phi\circ U)$. ∎
Proof of Proposition˜1.
By Lemma˜1, the conditions in (b) and (c) are sufficient for, respectively, property (1enumi) and property (2enumi). By Lemma˜3, the condition in (b) is necessary for property (1enumi), and by Lemma˜4, the condition in (c) is necessary for property (2enumi).
It remains to derive the characterisation of property (1) by that of property (1enumi) and to show that property (2enumi) is equivalent to property (2). To that end, note that by the above, property (1) holds if and only if
  1. (i)
    $\phi_{s}$ is weakly increasing for all $s\in\mathcal{S}$ and, for any $S\subseteq\mathcal{S}$ with $\mathopen{}\mathclose{{\left\lvert S}}\right\rvert\geq 2$, $(\phi_{s})_{s\in S}$ satisfies the conditions stated in Proposition˜1(b)
(where the ‘only if’ direction follows by applying property (1) to finite problems, $(X,S,U)$, for each non-empty subset $S\subseteq\mathcal{S}$ of situations), that property (2) holds if
  1. (ii)
    $\phi_{s}$ is weakly increasing for all $s\in\mathcal{S}$ and, for any $S\subseteq\mathcal{S}$ with $\mathopen{}\mathclose{{\left\lvert S}}\right\rvert\geq 2$, $(\phi_{s})_{s\in S}$ satisfies the conditions stated in Proposition˜1(c),
and that property (2) holds only if
  1. (iii)
    for any $S\subseteq\mathcal{S}$ with $\mathopen{}\mathclose{{\left\lvert S}}\right\rvert\geq 2$, $(\phi_{s})_{s\in S}$ satisfies the conditions stated in Proposition˜1(c).
This concludes the proof since (i) is equivalent to $(\phi_{s})_{s\in\mathcal{S}}$ satisfying the properties stated in Proposition˜1(a), and both (ii) and (iii) are equivalent to $(\phi_{s})_{s\in\mathcal{S}}$ satisfying the properties stated in Proposition˜1(c). ∎
Anhang B Proof of Proposition˜2
We establish the proposition by assembling a collection of lemmata.
Lemma 5.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. Properties (3) and (4) hold if for each $s\in\mathcal{S}$, $\phi_{s}$ is convex and strictly increasing.
Beweis.
Suppose that for every $s\in\mathcal{S}$, $\phi_{s}$ is convex and strictly increasing. Fix any $x\notin\text{NWD}(X,S,U)$ ($x\notin\text{NSD}(X,S,U)$); we will show that $x\notin\text{NWD}(X,S,\Phi\circ U)$ ($x\notin\text{NSD}(X,S,\Phi\circ U)$). By hypothesis, there is a mixture $p\in\Delta(X)$ such that
$\sum_{y\in X}U(y,\cdot)p(y)\underset{(\gg)}{>}U(x,\cdot).$
By Pratt’s theorem, it must, therefore, be that
$\sum_{y\in X}\Phi(U(y,\cdot))p(y)\underset{(\gg)}{>}\Phi(U(x,\cdot)),$
as desired.∎
Lemma 6.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. If either property (3enumi) or property (4enumi) holds, then for each $s\in\mathcal{S}$, $\phi_{s}:\mathbf{R}\to\mathbf{R}$ is strictly increasing and continuous.
Beweis.
Suppose that either property (3enumi) or property (4enumi) holds. Then, it is immediate that each $\phi_{s}$ must be strictly increasing: if for some $s\in\mathcal{S}$ and some $a<b$, $\phi_{s}(a)\geq\phi_{s}(b)$, it is easy to construct a problem where some $x$ is dominated (either weakly or strictly) pre-transformation and undominated after, due to a strict inequality in $s$ failing after the transformation.
For all $u\in\mathbf{R}$, let
$\phi_{s}(u-)=\sup_{(-\infty,u)}\phi_{s}\quad\text{and}\quad\phi_{s}(u+)=\inf_{(u,\infty)}\phi_{s}.$
Suppose that $\phi_{s}$ is discontinuous at $u\in\mathbf{R}$ for some $s\in\mathcal{S}$ and seek a contradiction. It suffices to exhibit $a,b,c\in\mathbf{R}^{\mathcal{S}}$ such that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$ and $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$, where $\mathrm{id}(x,\cdot)\coloneqq x$ for each $x\in\{a,b,c\}$. To this end, note that
$\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ (1)
for $v>u$ sufficiently close to $u$, since $\phi_{s}$ is strictly increasing and discontinuous at $u$. Fix $t\in\mathcal{S}\setminus\{s\}$ and note that $\phi_{t}$ has countably many discontinuities since it is strictly increasing. Then we may choose $v>u$ such that (1) holds and $\phi_{t}$ is continuous at $v$. Since $\phi_{t}$ is continuous at $v$ and strictly increasing, there exists $w\in(u,v)$ such that
$\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}.$ (2)
Given $v^{\prime}<u$ and $w^{\prime}\in(u,w)$, let $a,b,c\in\mathbf{R}^{\mathcal{S}}$ be given by
$a_{r}=\begin{cases}v&\text{if $r=s$}\\ v^{\prime}&\text{if $r\neq s$}\end{cases}\qquad b_{r}=\begin{cases}v^{\prime}&\text{if $r=s$}\\ v&\text{if $r\neq s$.}\end{cases}\qquad c_{r}=\begin{cases}w^{\prime}&\text{if $r=s$}\\ w&\text{if $r\neq s$.}\end{cases}$
Since $\max\{u,w\}<v$, we may choose $v^{\prime}$ and $w^{\prime}$ sufficiently close to $u$ to ensure that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$.
It remains to show that $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$. To this end, note that $b_{s}<c_{s}<a_{s}$ and $a_{t}<c_{t}<b_{t}$, so that $\Phi_{s}(b)<\Phi_{s}(c)<\Phi_{s}(a)$ and $\Phi_{t}(a)<\Phi_{t}(c)<\Phi_{t}(b)$ (since $\Phi$ is situation-wise and each $\phi_{r}$ is strictly increasing). Then, it suffices to show that
$\Phi_{r}(c)>\frac{\Phi_{r}(a)+\Phi_{r}(b)}{2}$
for each $r\in\{s,t\}$. For the case $r=s$, note that
$\Phi_{s}(c)=\phi_{s}(w^{\prime})\geq\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ $\geq\frac{\phi_{s}(v^{\prime})+\phi_{s}(v)}{2}$
$=\frac{\Phi_{s}(b)+\Phi_{s}(a)}{2},$
where the first inequality holds since $w^{\prime}>u$, the second is (1), and the third holds since $u>v^{\prime}$. For the case $r=t$, note that
$\Phi_{t}(c)=\phi_{t}(w)\geq\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}$ $\geq\frac{\phi_{t}(v^{\prime})+\phi_{t}(v)}{2}$
$=\frac{\Phi_{t}(a)+\Phi_{t}(b)}{2},$
where the first inequality holds by monotonicity, the second is (2), and the third holds since $u>v^{\prime}$.
Accordingly, $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)\subseteq\mathrm{NSD}(\{a,b,c\},\mathcal{S},\Phi)$ but $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})\supseteq\mathrm{NWD}(\{a,b,c\},\mathcal{S},\mathrm{id})$, contradicting properties (3enumi) and (4enumi). ∎
Lemma 7.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. Fix a family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$, and let $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. If either property (3enumi) or property (4enumi) holds, then for each $s\in\mathcal{S}$, $\phi_{s}:\mathbf{R}\to\mathbf{R}$ is strictly increasing and convex.
Beweis.
Suppose that either property (3enumi) or property (4enumi) holds. Then, by Lemma 6, each $\phi_{s}$ must be strictly increasing and continuous.
To establish convexity, suppose first that property (3enumi) holds. Since each $\phi_{s}$ is strictly increasing and continuous, each image $I_{s}\coloneqq\phi_{s}(\mathbf{R})$ is a (non-empty) open interval of $\mathbf{R}$, and $\phi_{s}$ admits an inverse $\phi_{s}^{-1}:I_{s}\to\mathbf{R}$. Since $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise, it admits an inverse $\Phi^{-1}:I\to\mathbf{R}^{\mathcal{S}}$, where $I\coloneqq\prod_{s\in\mathcal{S}}I_{s}$. Note that, given any finite problem $(X,\mathcal{S},V)$ with $V(X)\subseteq I$, $\mathrm{NSD}(X,\mathcal{S},V)\subseteq\mathrm{NSD}(X,\mathcal{S},\Phi^{-1}\circ V)$. Then, Lemma˜3 implies that each $\phi_{s}^{-1}$ is concave; thus, each $\phi_{s}$ is convex.
The same steps, with Lemma˜4 used in place of Lemma˜3, deliver convexity of each $\phi_{s}$ for the case in which property (4enumi) holds. ∎
Proof of Proposition˜2.
By Lemma˜5, if $\phi_{s}$ is convex and strictly increasing for each $s\in\mathcal{S}$, then properties (3) and (4) hold, which implies that properties (3enumi) and (4enumi) hold. The equivalence is closed by Lemma 7. ∎
Anhang C Proof of Theorem˜1
In light of Propositions˜1 and 2, what must be shown is that $\Phi$ being situation-wise is necessary for expansion and for contraction of NSD and of NWD. We do this via two lemmata.
Lemma 8.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$. If property (1) or property (2) holds, then $\Phi$ is situation-wise.
Proof of Lemma˜8.
Fix $s\in\mathcal{S}$ and take any $u,v\in\mathbf{R}^{\mathcal{S}}$ with $u_{s}=v_{s}$. Consider the finite problem $(\{x,y\},\{s\},U)$, where $x\neq y$, $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$. Since $u_{s}=v_{s}$, neither object weakly dominates the other, so $\{x,y\}=\text{NWD}(\{x,y\},\{s\},U)=\text{NSD}(\{x,y\},\{s\},U)$. Assuming that property (2) (property (1)) holds, $\{x,y\}=\text{NWD}(\{x,y\},\{s\},\Phi\circ U)$ ($\{x,y\}=\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$) as well. Then $\Phi_{s}(u)=\Phi_{s}(v)$, or else the object with the larger $s$-coordinate after the transformation strictly dominates the other on $S=\{s\}$. Since $\Phi_{s}(u)=\Phi_{s}(v)$ for all $u,v\in\mathbf{R}^{\mathcal{S}}$ with $u_{s}=v_{s}$, there exists a function $\phi_{s}:\mathbf{R}\to\mathbf{R}$ with $\Phi_{s}(w)=\phi_{s}(w_{s})$ for all $w$. Since $s$ was arbitrary, $\Phi$ is situation-wise. ∎
Lemma 9.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$, and consider a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$. If property (3) or property (4) holds, then there exists a family $(\phi_{s})_{s\in\mathcal{S}}$ of continuous and strictly increasing functions $\mathbf{R}\to\mathbf{R}$ such that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise.
Proof of Lemma˜9.
For each $s\in\mathcal{S}$, let $\pi^{s}:\mathbf{R}\to\mathbf{R}^{S}$ be given by $\pi^{s}_{s}(u)=u$ and $\pi^{s}_{t}(u)=0$ for all $t\neq s$, and set $\phi_{s}=\Phi_{s}\circ\pi^{s}$. Recall that for all $u\in\mathbf{R}$,
$\phi_{s}(u-)=\sup_{(-\infty,u)}\phi_{s}\quad\text{and}\quad\phi_{s}(u+)=\inf_{(u,\infty)}\phi_{s}.$
Note that, for all $a,b\in\mathbf{R}^{\mathcal{S}}$ such that $a_{s}>b_{s}$, $\Phi_{s}(a)>\Phi_{s}(b)$. This implies that $\phi_{s}$ is strictly increasing and that $\phi_{s}(a_{s}-\varepsilon)<\Phi_{s}(a)<\phi_{s}(a_{s}+\varepsilon)$ for all $a\in\mathbf{R}^{\mathcal{S}}$ and $\varepsilon>0$. Letting $\varepsilon$ vanish yields
$\phi_{s}(a_{s}-)\leq\Phi_{s}(a)\leq\phi_{s}(a_{s}+)\quad\text{for all $a\in\mathbf{R}^{\mathcal{S}}$.}$ (3)
Then, it is enough to show that $\phi_{s}$ is continuous for each $s\in\mathcal{S}$, as this implies that $\Phi_{s}(a)=\phi_{s}(a_{s})$ for all $a\in\mathbf{R}^{\mathcal{S}}$.
We now mimic the proof of Lemma˜6 nearly verbatim. Suppose that $\phi_{s}$ is discontinuous at $u\in\mathbf{R}$ for some $s\in\mathcal{S}$ and seek a contradiction. It suffices to exhibit $a,b,c\in\mathbf{R}^{\mathcal{S}}$ such that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$ and $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$, where $\mathrm{id}(x,\cdot)\coloneqq x$ for each $x\in\{a,b,c\}$. To this end, note that
$\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ (4)
for $v>u$ sufficiently close to $u$, since $\phi_{s}$ is strictly increasing and discontinuous at $u$. Fix $t\in\mathcal{S}\setminus\{s\}$ and note that $\phi_{t}$ has countably many discontinuities since it is strictly increasing. Then we may choose $v>u$ such that (4) holds and $\phi_{t}$ is continuous at $v$. Since $\phi_{t}$ is continuous at $v$ and strictly increasing, there exists $w\in(u,v)$ such that
$\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}.$ (5)
Given $v^{\prime}<u$ and $w^{\prime}\in(u,w)$, let $a,b,c\in\mathbf{R}^{\mathcal{S}}$ be given by
$a_{r}=\begin{cases}v&\text{if $r=s$}\\ v^{\prime}&\text{if $r\neq s$}\end{cases}\qquad b_{r}=\begin{cases}v^{\prime}&\text{if $r=s$}\\ v&\text{if $r\neq s$.}\end{cases}\qquad c_{r}=\begin{cases}w^{\prime}&\text{if $r=s$}\\ w&\text{if $r\neq s$.}\end{cases}$
Since $\max\{u,w\}<v$, we may choose $v^{\prime}$ and $w^{\prime}$ sufficiently close to $u$ to ensure that $c\notin\mathrm{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$.
It remains to show that $c\in\mathrm{NWD}(\{a,b,c\},\mathcal{S},\Phi)$. To this end, note that $b_{s}<c_{s}<a_{s}$ and $a_{t}<c_{t}<b_{t}$, so that $\Phi_{s}(b)<\Phi_{s}(c)<\Phi_{s}(a)$ and $\Phi_{t}(a)<\Phi_{t}(c)<\Phi_{t}(b)$, from above. Then, it suffices to show that
$\Phi_{r}(c)>\frac{\Phi_{r}(a)+\Phi_{r}(b)}{2}$
for each $r\in\{s,t\}$. For the case $r=s$, note that
$\Phi_{s}(c)\geq\phi_{s}(w^{\prime}-)\geq\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ $\geq\frac{\phi_{s}(v^{\prime})+\phi_{s}(v)}{2}$
$=\frac{\Phi_{s}(b)+\Phi_{s}(a)}{2},$
where the first inequality and the equality follow from (3), the second inequality holds since $w^{\prime}>u$, the third is (4), and the fourth holds since $u>v^{\prime}$. For the case $r=t$, note that
@@
where the first inequality and the equality follow from (3), the second inequality is (5), and the third holds since $u>v^{\prime}$. ∎
Proof of Theorem˜1.
Propositions˜1 and 2 establish sufficiency direction. For necessity, in light of Propositions˜1 and 2, it suffices to show that $\Phi$ being situation-wise is necessary for each of properties (1)(4), and that follows from Lemmata˜9 and 8. ∎
Anhang D Proof of Theorem˜2
@@
The ‘if’ parts of Theorem˜2(a)(b) are easily established.
For the ‘only if’ parts, mimicking Theorem˜1’s proof exactly, we conclude that $\Phi$ being $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise (for some family $(\phi_{s})_{s\in\mathcal{S}}$ of maps $\mathbf{R}\to\mathbf{R}$) is necessary for both properties ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) to hold. We now argue that $\phi_{s}$ being weakly increasing for each $s\in\mathcal{S}$ is necessary as well. Suppose that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise for some family $(\phi_{s})_{s\in\mathcal{S}}$, but there are $s\in\mathcal{S}$ and $u_{s}<v_{s}$ in $\mathbf{R}$ such that $\phi_{s}(u_{s})>\phi_{s}(v_{s})$, and consider a finite problem $(\{x,y\},\mathcal{S},U)$ with $U$ such that $U(x,s)=u_{s}$ and $U(y,s)=v_{s}$. Note that
$\widehat{\text{NSD}}(X,\{s\},U)=\{y\}=\widehat{\text{NWD}}(X,\{s\},U)$
and that
$\widehat{\text{NSD}}(X,\{s\},\Phi\circ U)=\{x\}=\widehat{\text{NWD}}(X,\{s\},\Phi\circ U),$
so that properties ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) both fail.

### v2

@@ 6 Dominance by objects (rather than mixtures)
[…]
Anhang G Proof of Theorem˜2
@@
The sufficiency (‘if’) claims in Theorem˜2(a)(d) are easily established. It remains to show necessity (the ‘only if’ claims). Fix arbitrary $x\neq y$.
For the necessity (‘only if’) claims in Theorem˜2(a)(b), assume that either property ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) or property ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) holds. Mimicking the proof of Lemma˜3(i) in appendix˜F yields that $\Phi$ must be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise for some family $(\phi_{s})_{s\in\mathcal{S}}$ of functions $\mathbf{R}\to\mathbf{R}$. Furthermore, $\phi_{s}$ must be weakly increasing for every $s\in\mathcal{S}$. To see why, suppose toward a contradiction that there are $s\in\mathcal{S}$ and $u_{s}>v_{s}$ in $\mathbf{R}$ such that $\phi_{s}(u_{s})<\phi_{s}(v_{s})$. Then for $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ satisfying $U(x,s)=u_{s}$ and $U(y,s)=v_{s}$, we have $x\in\widehat{\text{NWD}}(\{x,y\},\{s\},U)\subseteq\widehat{\text{NSD}}(\{x,y\},\{s\},U)$ and $x\notin\widehat{\text{NSD}}(\{x,y\},\{s\},\Phi\circ U)\supseteq\widehat{\text{NWD}}(\{x,y\},\{s\},\Phi\circ U)$, so that properties ($\hskip 1.00006pt\widehat{1}\hskip 1.00006pt$) and ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) both fail—a contradiction.
This establishes the ‘only if’ part of Theorem˜2(a). To establish the ‘only if’ part of Theorem˜2(b), suppose that there is an $s\in\mathcal{S}$ such that $\phi_{s}$ is not strictly increasing and that there is at least one $t\in\mathcal{S}$ such that $\phi_{t}$ is non-constant; we must show that property ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) fails.
  • Suppose first that there is a $t\in\mathcal{S}\setminus\{s\}$ such that $\phi_{t}$ is non-constant. Then (recalling that $\phi_{s}$ and $\phi_{t}$ are weakly increasing) there are $u>v$ and $u^{\prime}>v^{\prime}$ in $\mathbf{R}$ such that $\phi_{s}(u)=\phi_{s}(v)$ and $\phi_{t}(u^{\prime})>\phi_{t}(v^{\prime})$. Hence for $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ satisfying $U(x,s)=u$, $U(x,t)=v^{\prime}$, $U(y,s)=v$ and $U(y,t)=u^{\prime}$, we have $\widehat{\text{NWD}}(\{x,y\},\{s,t\},U)\ni x\notin\widehat{\text{NWD}}(\{x,y\},\{s,t\},\Phi\circ U)$, so property ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) fails.
  • Suppose instead that $\phi_{t}$ is constant for every $t\in\mathcal{S}\setminus\{s\}$. Then $\phi_{s}$ is (weakly increasing and) non-constant, so there are $u>v$ such that $\phi_{s}(u)>\phi_{s}(v)$. Fix arbitrary $t\in\mathcal{S}\setminus\{s\}$ and $u^{\prime}>v^{\prime}$ in $\mathbf{R}$. Then for $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ satisfying $U(x,s)=v$, $U(x,t)=u^{\prime}$, $U(y,s)=u$ and $U(y,t)=v^{\prime}$, we have $\widehat{\text{NWD}}(\{x,y\},\{s,t\},U)\ni x\notin\widehat{\text{NWD}}(\{x,y\},\{s,t\},\Phi\circ U)$, so property ($\hskip 1.00006pt\widehat{2}\hskip 1.00006pt$) fails.
For the necessity (‘only if’) claims in Theorem˜2(c)(d), assume that either property ($\hskip 1.00006pt\widehat{3}\hskip 1.00006pt$) or property ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$) holds. Mimicking (the beginning of) the proof of Lemma˜3(ii) in appendix˜F yields that $\Phi$ must strictly situation-wise monotone in the sense that for any $u,v\in\mathbf{R}^{\mathcal{S}}$ and any situation $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$.
This establishes the ‘only if’ part of Theorem˜2(c). To establish the ‘only if’ part of Theorem˜2(d), suppose that there are $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s,t\in\mathcal{S}$ such that $u_{s}>v_{s}$, $u_{r}\geq v_{r}$ for every $r\in\mathcal{S}\setminus\{s\}$, and $\Phi_{t}(u)<\Phi_{t}(v)$; we must show that property ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$) fails. By strict situation-wise monotonicity, $\Phi_{s}(u)>\Phi_{s}(v)$ and (thus) $s\neq t$. Then for $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ given by $U(x,\cdot)\coloneqq v$ and $U(y,\cdot)\coloneqq u$, we have $\widehat{\text{NWD}}(\{x,y\},\{s,t\},\Phi\circ U)\ni x\notin\widehat{\text{NWD}}(\{x,y\},\{s,t\},U)$, so property ($\hskip 1.00006pt\widehat{4}\hskip 1.00006pt$) fails. ∎
Anhang H Proof of the claim in Remark˜2
We will exhibit a set $\mathcal{S}$ with $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$ and a map $\Phi:\mathbf{Z}^{\mathcal{S}}\to\mathbf{Z}^{\mathcal{S}}$ that is not situation-wise, and yet satisfies properties (3) and (4) when restricted to finite problems with integer-valued payoffs, i.e. satisfies the properties
  1. ($3_{\mathbf{Z}}$)
    $\text{NSD}(X,S,\Phi\circ U)\subseteq\text{NSD}(X,S,U)$ for every finite problem $(X,S,U)$ such that $U(x,\cdot)\in\mathbf{Z}^{\mathcal{S}}$ for each $x\in X$.
  2. ($4_{\mathbf{Z}}$)
    $\text{NWD}(X,S,\Phi\circ U)\subseteq\text{NWD}(X,S,U)$ for every finite problem $(X,S,U)$ such that $U(x,\cdot)\in\mathbf{Z}^{\mathcal{S}}$ for each $x\in X$.
To that end, let $\mathcal{S}\coloneqq\{1,2\}$, and define $\Phi:\mathbf{Z}^{2}\to\mathbf{Z}^{2}$ by